# OpticalRS-4M: Scaling Efficient Masked Autoencoder Learning on Large Remote Sensing Dataset

### v3

@@ Abstract
Masked Image Modeling (MIM) has become an essential method for building foundational visual models in remote sensing (RS). However, the limitations in size and diversity of existing RS datasets restrict the ability of MIM methods to learn generalizable representations. Additionally, conventional MIM techniques, which require reconstructing all tokens, introduce unnecessary computational overhead. To address these issues, we present a new pre-training pipeline for RS models, featuring the creation of a large-scale RS dataset and an efficient MIM approach. We curated a high-quality dataset named OpticalRS-4M by collecting publicly available RS datasets and processing them through exclusion, slicing, and deduplication. OpticalRS-4M comprises 4 million optical images covering various RS tasks, such as object detection and pixel segmentation. To enhance efficiency, we propose SelectiveMAE, a pre-training method that dynamically encodes and reconstructs semantically rich patch tokens, thereby reducing the inefficiencies of traditional MIM models caused by redundant background pixels in RS images. Extensive experiments demonstrate that OpticalRS-4M significantly improves classification, detection, and segmentation performance, while SelectiveMAE increases training efficiency over 2$\times$ times. This highlights the effectiveness and scalability of our pipeline in developing RS foundational models. The dataset, source code, and trained models will be released at OpticalRS-4M.
@@ 1 Introduction
To address these challenges, we propose a new pipeline to collect, create, and efficiently process a large optical RS dataset. Firstly, we reviewed publicly available remote sensing datasets from the past decade and selected them based on the DiRS (Diversity, Richness, and Scalability) principle [20]. Nevertheless, issues such as inconsistent data sources, excessively large image sizes, and redundant pixels still exist. Therefore, we also apply exclusion, slicing, and deduplication processes to further improve data quality. Correspondingly, we obtain a large-scale RS dataset called OpticalRS-4M comprising 4 million optical images, which is designed to fully leverage the representation learning capabilities of MIM methods in RS applications. OpticalRS-4M exceeds previous RS datasets [24, 21, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 18, 37, 38], being at least four times larger. Moreover, OpticalRS-4M encompasses a wide range of diverse RS scenarios encountered in downstream tasks such as object-level detection and pixel-level segmentation (Fig. 1 left).
@@
Despite substantial efforts in training RSFM using MIM methods, such as leveraging general image knowledge [25, 26, 28], expanding parameter scales [29], integrating spatio-temporal information [30, 31, 32], and learning multi-scale features [33, 35, 36], the computational burden and slow convergence when employing MIM training on large-scale RS datasets cannot be ignored. Specifically, pre-training on 1 million RS samples requires 107 hours for the ViT-B [39] backbone on 8 Nvidia A100 GPUs [18]. This issue becomes even more pronounced when training on larger datasets, e.g., OpticalRS-4M. In natural scene analysis, this issue has led to numerous studies [40, 41, 42, 43, 44, 45] aimed at improving MIM training efficiency. One approach is to accelerate the token reconstruction process by using decoders with fewer parameters [40, 41]. Another approach is to reduce the number of visible patch tokens input into the vision encoder [42, 43, 44], speeding up feature extraction.
@@
However, conventional MIM approaches, such as the encoding-then-decoding procedure, overlook the unique characteristics of RS images, which typically feature sparse foreground pixels and dense backgrounds [18, 28]. This raises two key questions about how to efficiently conduct MIM training in the RS field: 1) Is it necessary to reconstruct all the redundant background patches during the MIM decoding process? 2) Is there a feasible way to encode fewer image patches ($\bm{e.g.,\leq}$25%) to accelerate the convergence of MIM training? To address the first question, a measure-based selection process is needed to identify the appropriate patches for reconstruction. For the second question, the intuition is that the patch tokens used in the encoding-then-decoding procedure should effectively capture feature dependencies in RS images.
@@
Regarding the above issues, in the second part of the pipeline, we introduce a MIM method called SelectiveMAE for efficiently processing RS images, which dynamically encodes and reconstructs patch tokens based on their semantic richness. Specifically, SelectiveMAE utilizes the Histogram of Oriented Gradients (HOG) algorithm to quantify the semantic richness of patches. Then, it selects a subset of patch tokens ($e.g.,\leq$50%) with higher HOG values for feature encoding ($e.g.,\leq$15%) and pixel reconstruction ($e.g.,\leq$35%). However, using an extremely low ratio of visible patches during MIM training can lead to gradient explosion or vanishing. To mitigate this, we designed a Progressive Semantic Token Selection (PSTS) module, which dynamically selects semantically relevant patch tokens during the entire training phase. In the beginning, SelectiveMAE encodes semantically rich tokens and reconstructs semantically similar ones to warm up the training process. As training advances, SelectiveMAE shifts to reconstructing high-semantic tokens from encoded lower-semantic ones to capture complementary semantic dependencies. This analogical-to-complementary strategy allows SelectiveMAE to efficiently and progressively learn robust representations of RS images while accelerating MIM convergence. Our experiments reveal that 40% of RS image patches are sufficient to train a comparable model, offering new insights into MIM training on RS images.
The contributions of this paper are summarized as follows:
@@
2) Using this pipeline, we create the OpticalRS-4M dataset, which is a large-scale RS dataset comprising 4 million optical images with diversified coverage scenarios.
@@
Experiment results suggest the effectiveness and scalability of the proposed pipeline. OpticalRS-4M significantly enhances the performance of RS foundation models in downstream tasks, while SelectiveMAE achieves over 2×2$\times$2 × speedup in pre-training compared to MAE [10].
@@ 2 Related Work
Remote Sensing Datasets. In recent years, many RS datasets have been created for tasks such as scene classification [23, 46], object detection [47, 48, 49], and segmentation [50, 51, 52]. The availability of free, unlabeled satellite images has led to the development of large-scale RS datasets. Some works combine various sensor data to create extensive datasets. SEN12MS [53] with 180,662 triplets of dual-pol synthetic aperture radar (SAR) and multi-spectral Sentinel-2 image patches. SSL4EO-S12 [54] dataset uses publicly available SAR/optical satellite data collected by the European Space Agency’s Sentinel mission. SSL4EO-L [55] consists entirely of multispectral data. SatlasPretrain [56] is constructed from data sources such as Sentinel-2, Sentinel-1, Landsat 8/9, and 0.5 - 2 meter/pixel aerial imagery. While it includes millions of images, the majority are multispectral and SAR, with only a small fraction being RGB. There are also some large-scale RS visual-language datasets such as Skyscript [57] and RS5M [58], but they primarily focus on multimodal tasks. Currently, some datasets focus solely on optical RS images. MillionAID [20] offers a million-scale dataset for RS optical image classification, while SeCo [22] and CACo [21] provide nearly a million images of the same location over different times. However, these datasets primarily target scene classification and often overlook fine-grained target information, limiting their utility for various downstream tasks. To address this gap, with the proposed pipeline, we introduce the OpticalRS-4M dataset, which is larger and more diverse, enhancing performance across multiple downstream tasks.
@@
Masked Image Modeling. Inspired by the success of Masked Language Modeling (MLM) in NLP [59], MIM has been developed for visual pre-training [60, 11, 10, 61, 62]. MIM learns image representations by reconstructing masked tokens, focusing on various regression targets [63, 64, 65, 66, 67], masking strategies [68, 69], and reconstruction methods [70, 71, 72, 73]. For instance, MAE [10] demonstrates that predicting pixel values can be as effective as using complex targets. A major challenge for MIM is its high computational demand and lengthy pre-training times. To mitigate this, some studies use asymmetric encoder-decoder strategies [40, 41], reduce input patches [42, 44] or use a novel Difficulty-Flatten Loss [43]. Additionally, CrossMAE [45] employs cross-attention between masked and visible tokens to enhance efficiency without sacrificing performance. However, these methods do not account for the unique characteristics of RS images, such as sparse foreground information and complex backgrounds. Considering these issues, we introduce an efficient MIM method named SelectiveMAE that significantly speeds up pre-training, enhancing the practicality of developing RS foundation models based on large-scale datasets.
@@
Remote Sensing Foundation Models. Despite the abundance of RS data, much of it remains unlabeled and thus inaccessible for supervised learning [74]. Self-supervised learning methods have recently been employed to extract representations from unlabeled RS data. Although there are some contrastive self-supervised methods [24, 21, 22, 75], recent advancements have primarily centered around generative self-supervised methods. For example, RVSA [18] leverages the vision transformer backbone pre-trained by the vanilla MAE [10] method and introduces computationally efficient rotated varied-size window attention to replace the original attention during fine-tuning. In addition, many studies are focusing on improving generative self-supervised algorithms by leveraging general image knowledge [25, 26, 28], scaling up parameter sizes [29], integrating spatio-temporal data [30, 31, 32], handling multi-sensor data [76, 77, 78, 79, 80], and employing multi-scale concepts [33, 35, 36]. However, these methods have not effectively addressed the substantial computational burden associated with self-supervised pre-training in RS. In the paper, we propose a new pipeline that can collect, create, and efficiently process large amounts of optical RS data for developing RS foundation models, enhancing the practicality of MIM pre-training on large-scale datasets.
3 Methodology
@@
3.1 Dataset Generation
Figure 1: Left: The organization of data sources in OpticalRS-4M. Right: The comparison between OpticalRS-4M and other datasets.
@@ 3.1 Dataset Generation
Recent progress in self-supervised pre-training for RSFMs is limited by the relatively small scale and diversity of existing RS datasets compared to natural scene datasets. To overcome this, we first curate a large-scale RS dataset, OpticalRS-4M, through extensive data collection and preprocessing.
3.1.1 Data Collection
@@
We began by reviewing publicly available RS datasets from the past decade. Recognizing that not all data are suitable for self-supervised pre-training, we adopted specific criteria for data collection, as outlined in DiRS [20]: 1) Diversity: A dataset is diverse if its images capture various typical visual features of relevant scenes or targets and offer complementarity. Intra-class Diversity: High diversity within classes ensures comprehensive representation of real-world target distributions. To ensure this, we consider the data from different sources of the same scene or target. Moreover, in the RS field, the attributes of objects may vary with geographic location and imaging time, so the change detection data is also incorporated. Inter-class Diversity: To enable self-supervised pre-training to distinguish between categories effectively, we incorporated as many fine-grained categories as possible, especially those with high semantic overlap. This strategy helps the model learn robust feature representations. 2) Richness: Beyond diversity, richness is also essential. We ensure diverse content features and large sample sizes by collecting images under various conditions — weather, seasons, lighting, and so on — introducing variability in translation, viewpoint, object posture, spatial resolution, background, and occlusion. Our data collection process actively includes the sample of the same category under different imaging conditions to ensure comprehensive real-world representation, enhancing the model’s representation and generalization capabilities. 3) Scalability: Given the evolving applications of remote sensing images, scalability is also crucial. We constantly enrich the dataset based on scene and target categories across classification, detection, and segmentation tasks.
In summary, following the guidelines, we constructed a large-scale RS image dataset with diverse coverage scenarios. Fig. 1 left illustrates the included RS datasets collected.
3.1.2 Data Preprocessing
@@
Following data collection, challenges such as inconsistent data sources, oversized images, and redundant pixels remained. To address these, we implemented a standardized, scalable data preprocessing workflow consisting of four steps:
@@
1) We focused exclusively on optical images for this study, excluding multispectral and SAR data, though future updates will incorporate these modalities for multi-modal self-supervised pre-training. 2) To manage the large image sizes, we randomly cropped high-resolution images into smaller slices. 3) We filtered out images with fewer than 64 or more than 1024 pixels. 4) The remaining images were combined and duplicates were removed using a two-phase approach: first, a coarse phase with perceptual hashing [81], and then a refined phase involving manual review. This ensured that only highly similar images were excluded. As a result of these steps, we created the OpticalRS-4M dataset, which contains approximately 4 million high-quality RS images — significantly larger than previous representative RS datasets, as illustrated in Fig. 1 right.
In summary, we followed the DiRS principles for data collection, applying exclusion, slicing, and deduplication to ensure high-quality results. We’ve screened and processed most small and medium-scale datasets from the past decade in this field. More efforts can be made to extend OpticalRS-4M by including larger datasets like SSL4EO-S12 [54] and SSL4EO-L [55] to enhance dataset capacity.
3.1.3 Dataset Diversity
Refer to captionRefer to caption
Figure 2: Left: Main categories with visual examples from OpticalRS-4M.Right: Full hierarchy of OpticalRS-4M categories.
@@ 3.1.3 Dataset Diversity
To demonstrate the broad scope of scenes and target categories covered by our dataset, we conducted a statistical analysis accompanied by a detailed label explanation, reviewing all the scenarios and targets included in OpticalRS-4M, as shown in Fig. 2. The analysis reveals that OpticalRS-4M comprises 12 main categories, each containing numerous subclasses. Unlike existing self-supervised optical remote sensing datasets, OpticalRS-4M not only offers more comprehensive scene and target information but also introduces “Events” as a distinct high-level category. This category includes labels such as “Concert”, “Parade/Protest”, “Religious Activity”, “Fire”, “Flood”, “Landslide”, “Post-Earthquake”, and “Traffic Collision”. This expanded information enhances the versatility of OpticalRS-4M for a wider range of downstream tasks.

### v4

# Scaling Efficient Masked Image Modeling on Large Remote Sensing Dataset

@@ Abstract
Masked Image Modeling (MIM) has become an essential method for building foundational visual models in remote sensing (RS). However, the limitations in size and diversity of existing RS datasets restrict the ability of MIM methods to learn generalizable representations. Additionally, conventional MIM techniques, which require reconstructing all tokens, introduce unnecessary computational overhead. To address these issues, we present a new pre-training pipeline for RS models, featuring the creation of a large-scale RS dataset and an efficient MIM approach. We curated a high-quality dataset named OpticalRS-13M by collecting publicly available RS datasets and processing them through exclusion, slicing, and deduplication. OpticalRS-13M comprises 13 million optical images covering various RS tasks, such as object detection and pixel segmentation. To enhance efficiency, we propose SelectiveMAE, a pre-training method that dynamically encodes and reconstructs semantically rich patch tokens, thereby reducing the inefficiencies of traditional MIM models caused by redundant background pixels in RS images. Extensive experiments demonstrate that OpticalRS-13M significantly improves classification, detection, and segmentation performance, while SelectiveMAE increases training efficiency over 2$\times$ times. This highlights the effectiveness and scalability of our pipeline in developing RS foundational models.
@@ 1 Introduction
Figure 1: Advantages of SelectiveMAE on finetuning performance and pre-training efficiency. The accuracy is evaluated on the DIOR detection dataset [41], while the training efficiency is measured on NVIDIA A100 GPUs, with different versions of ViT [19].
@@
Figure 2: Advantages of OpticalRS-13M in data volumn and diversity. (a) The volumn of OpticalRS-13M surpasses at least 4$\times$ compared to existing RS datasets. (b) OpticalRS-13M includes significantly more sample categories than others, spanning both object and event types. (c) OpticalRS-13M provides more abundant feature representations compared to other large-scale RS datasets. Here, 1,000 samples are separately selected from different datasets, and features are extracted using a pre-trained ViT-B.
@@
To address these challenges, taking an example of visible light RS imagery, we propose a new pipeline to collect, create, and efficiently process a large RS dataset. Firstly, we reviewed publicly available remote sensing datasets from the past decade and selected them based on the DiRS (Diversity, Richness, and Scalability) principle [48]. Nevertheless, issues such as inconsistent data sources, excessively large image sizes, and redundant pixels still exist. Therefore, we also apply exclusion, slicing, and deduplication processes to further improve data quality. Correspondingly, we obtain a large-scale RS dataset called OpticalRS-13M comprising 13 million images, which is designed to fully leverage the representation learning capabilities of MIM methods in RS applications. OpticalRS-13M exceeds previous RS datasets [1, 50, 51, 33, 71, 65, 10, 17, 92, 35, 42, 56, 72, 64, 53, 70, 63, 52], being at least four times larger (Fig. 2 left). Moreover, OpticalRS-13M encompasses a wide range of diverse RS scenarios encountered in downstream tasks such as object-level detection and pixel-level segmentation (Fig. 2 right).
@@
Despite substantial efforts in training RSFM using MIM methods [51, 18, 17, 56], the computational burden and slow convergence when employing MIM training on large-scale RS datasets cannot be ignored. Specifically, pre-training on 1 million RS samples requires 107 hours for the ViT-B [19] backbone on 8 Nvidia A100 GPUs [70]. This issue becomes even more pronounced when training on larger datasets, e.g., OpticalRS-13M. In natural scene analysis, this issue has led to numerous studies [25, 75, 40, 11, 23, 20] aimed at improving MIM training efficiency. One approach is to accelerate the token reconstruction process by using decoders with fewer parameters [25, 75]. Another approach is to reduce the number of visible patch tokens input into the vision encoder [40, 11, 23], speeding up feature extraction.
@@
However, conventional MIM approaches, such as the encoding-then-decoding procedure, overlook the unique characteristics of RS images, which typically feature sparse foreground pixels and dense backgrounds [70, 65]. This raises two key questions about how to efficiently conduct MIM training in the RS field: 1) Is it necessary to reconstruct all the redundant background patches during the MIM decoding process? 2) Is there a feasible way to encode fewer image patches (e.g., $\leq$25%) to accelerate the convergence of MIM training? To address the first question, a measure-based selection process is needed to identify the appropriate patches for reconstruction. For the second question, the intuition is that the patch tokens used in the encoding-then-decoding procedure should effectively capture feature dependencies in RS images.
@@
Regarding the above issues, in the second part of the pipeline, we introduce a MIM method called SelectiveMAE for efficiently processing RS images, which dynamically encodes and reconstructs patch tokens based on their semantic richness. Specifically, SelectiveMAE utilizes the Histogram of Oriented Gradients (HOG) algorithm to quantify the semantic richness of patches. Then, it selects a subset of patch tokens (e.g., $\leq$50%) with higher HOG values for feature encoding (e.g., $\leq$15%) and pixel reconstruction (e.g., $\leq$35%). However, using an extremely low ratio of visible patches during MIM training can lead to gradient explosion. To mitigate this, we designed a Progressive Semantic Token Selection (PSTS) module, which dynamically selects semantically relevant patch tokens during the entire training phase. In the beginning, SelectiveMAE encodes semantically rich tokens and reconstructs semantically similar ones to warm up the training process. As training advances, SelectiveMAE shifts to reconstructing high-semantic tokens from encoded lower-semantic ones to capture complementary semantic dependencies. This analogical-to-complementary strategy allows SelectiveMAE to efficiently and progressively learn robust representations of RS images while accelerating MIM convergence (Fig. 1). Our experiments reveal that 40% of RS image patches are sufficient to train a comparable model, offering new insights into MIM training on RS images.
In summary, our main contributions are as follows:
@@
2) Using this pipeline, we create the OpticalRS-13M dataset, which is a large-scale RS dataset comprising 13 million optical images with diversified coverage scenarios.
@@
Experiment results suggest the effectiveness and scalability of the proposed pipeline. OpticalRS-13M significantly enhances the performance of RS foundation models in downstream tasks, while SelectiveMAE achieves over 2×2$\times$2 × speedup in pre-training compared to MAE [28].
@@ 2 Related Work
Remote Sensing Datasets. In recent years, many RS datasets have been created for tasks such as scene classification [62, 84], object detection [16, 21, 38], and segmentation [2, 76, 68]. The availability of free, unlabeled satellite images has led to the development of large-scale RS datasets. Some works combine various sensor data to create extensive datasets. The SEN12MS dataset [59] consists of 180,662 triplets, each containing synthetic aperture radar (SAR) and multispectral Sentinel-2 image patches, and MODIS land cover maps. While SSL4EO-S12 [79] , SSL4EO-L [61] and SatlasPretrain [5] contain millions of images, the major data type in these datasets are multispectral and SAR, with only a small fraction being RGB. There are also some large-scale RS visual-language datasets such as Skyscript [80] and RS5M [96], but they primarily focus on multimodal tasks. Currently, there are also some datasets that focus solely on visible light RS images. MillionAID [48], SeCo [49] and CACo [50] provide nearly a million images of the same location over different times. However, these datasets primarily target scene classification and often overlook fine-grained target information, limiting their utility for various downstream tasks. To address this gap, with the proposed pipeline, we introduce the OpticalRS-13M dataset, which is larger and more diverse, enhancing performance across multiple downstream tasks.
@@
Masked Image Modeling. Inspired by the success of Masked Language Modeling (MLM) in NLP [9], MIM has been developed for visual pre-training [12, 89, 28, 66, 37]. MIM learns image representations by reconstructing masked tokens, focusing on various regression targets [29, 4, 82, 97, 47], masking strategies [60, 39], and reconstruction methods [67, 83, 45, 36]. A major challenge for MIM is its high computational demand and lengthy pre-training times. To mitigate this, approaches such as asymmetric encoder-decoder strategies [25, 75], reducing input patches [40, 23], or the Difficulty-Flatten Loss [11] have been proposed. Additionally, CrossMAE [20] employs cross-attention between masked and visible tokens to enhance efficiency without sacrificing performance. However, these methods do not account for the unique characteristics of RS images, such as sparse foreground information and complex backgrounds. Considering these issues, we introduce an efficient MIM method named SelectiveMAE that significantly speeds up pre-training, enhancing the practicality of developing RS foundation models based on large-scale datasets.
@@
Remote Sensing Foundation Models. Despite the abundance of RS data, much of it remains unlabeled and thus inaccessible for supervised learning [69]. Self-supervised learning methods have recently been employed to extract representations from unlabeled RS data. Due to the inefficiency of designing pretext tasks and gathering required data for contrastive self-supervised methods [1, 50, 49, 81], recent advancements have primarily centered around generative self-supervised methods, especially in MIM. Specifically, many studies aim to improve generative self-supervised algorithms by leveraging general image knowledge [51, 33, 65], scaling up parameter sizes [10, 18], integrating spatio-temporal information [17, 92, 42], encompassing geometric attributes [70, 46], handling multi-sensor data [34, 24, 27, 8, 31], and employing multi-scale concepts [56, 64, 53]. However, these methods have not effectively addressed the substantial computational burden associated with self-supervised pre-training in RS. In the paper, we propose a new pipeline that can collect, create, and efficiently process large amounts of optical RS data for developing RS foundation models, enhancing the practicality of MIM pre-training on large-scale datasets.
3 Method
@@
3.1 Dataset Curation
@@
Recent progress in self-supervised pre-training for RSFMs is limited by the relatively small scale and diversity of existing RS datasets compared to natural scene datasets. To overcome this, we curate a large-scale RS dataset, OpticalRS-13M, through extensive data collection and preprocessing.
@@
Data Collection We began by reviewing publicly available RS datasets from the past decade. Recognizing that not all data are suitable for self-supervised pre-training, we adopted specific criteria for data collection, as outlined in DiRS [48]: 1) Diversity: A dataset is diverse if its images capture various typical visual features of relevant scenes or targets and offer complementarity. Intra-class Diversity: High within-class diversity ensures comprehensive representation of real-world target distributions. We include data from different sources of the same scene or target, incorporating change detection data to capture object variations by geographic location and imaging time. Inter-class Diversity: To enable self-supervised pretraining to distinguish categories effectively, we included as many fine-grained categories as possible, especially those with high semantic overlap, helping the model learn robust feature representations. 2) Richness: Beyond diversity, richness is essential. We ensure varied content features and large sample sizes by collecting images under multiple conditions (e.g., weather, seasons, lighting), introducing variability in translation, viewpoint, object posture, spatial resolution, background, and occlusion. Sampling across imaging conditions provides a comprehensive real-world representation, enhancing model’s representation and generalization capabilities. 3) Scalability: Scalability is crucial given the evolving applications of remote sensing images. We continuously expand the dataset based on scene and target categories across classification, detection, and segmentation tasks. In summary, following the guidelines, we constructed a large-scale RS image dataset with diverse coverage scenarios.
@@
Data Preprocessing Following data collection, challenges such as inconsistent data sources, oversized images, and redundant pixels remained. To address these, we implemented a standardized, scalable data preprocessing workflow consisting of four steps:
@@
1) We focused exclusively on visible light images for this study, excluding multispectral and SAR data, though future updates will incorporate these modalities for multi-modal self-supervised pre-training. 2) To manage the large image sizes, we randomly cropped high-resolution images into smaller slices. 3) We filtered out images with fewer than 64 or more than 1024 pixels. 4) The remaining images were combined and duplicates were removed using a two-phase approach: first, a coarse phase with perceptual hashing [55], and then a refined phase involving manual review. This ensured that only highly similar images were excluded. As a result of these steps, we created the OpticalRS-13M dataset, which contains approximately 13 million high-quality RS images, significantly larger than previous representative RS datasets, as illustrated in Fig. 2 left. In summary, we followed the DiRS principles for data collection, applying exclusion, slicing, and deduplication to ensure high-quality. We’ve screened and processed most small and medium-scale datasets from the past decade in this field.
@@
Dataset Diversity To demonstrate the broad scope of scenes and target categories covered by our dataset, we conducted a statistical analysis accompanied by a detailed label explanation, as shown in Fig. 2 right. The analysis reveals that OpticalRS-13M comprises 12 main categories, each containing numerous subclasses. Unlike existing self-supervised optical remote sensing datasets, OpticalRS-13M not only offers more comprehensive scene and target information but also introduces “Events” as a distinct high-level category. This category includes labels such as “Fire”, “Flood”, “Landslide”, “Post-Earthquake”, and so on. This expanded information enhances the versatility of OpticalRS-13M for a wider range of downstream tasks.